# The Streaming Batch Model for Efficient and Fault-Tolerant Heterogeneous Execution

### v4

@@ 4.3.2 Input Rate Control
1:$P\leftarrow 0$triangleright Total processing time per partition
2:$\alpha_{0}\leftarrow 1$triangleright $\alpha_{i}:=$ Input:Output size ratio for $op_{i}$
3:for $i\leftarrow 1$ to $numOps$ do
@@
14:$budget\leftarrow budget+\mathrm{outputPartitionSize}(source)/P$
@@
  • Consider the first non-source operator: transform. Assume that the number of available execution slots to run the task is $E_{1}=6$ (out of 8 CPU slots). Assume the average task duration is $T_{1}=12$ seconds. Then the processing time of this stage is $P_{1}=T_{1}/E_{1}\cdot\alpha_{0}=12/6\times 1=2$, where $\alpha_{0}$, the output multiplier, is initialized to 1. In other words, transform takes 2s to process a source partition on average.
@@
If the run-time estimates of each operator’s processing rates are perfectly accurate, i.e. if there is no variance in the processing rates, it can be shown that the schedule produced is optimal. However, when there is variance in processing times or output sizes, the budget algorithm could overestimate the overall processing rate. Nevertheless, the algorithm is stable because it creates a negative feedback loop. If it overestimates the pipeline processing rate, more source tasks might launch and temporarily cause backpressure, or objects in the buffer to spill to disk. However, since these tasks still occupy execution slots, they will reduce the parallelism of downstream operators and lower the replenishment rate of the budget, which in turn limits the source task launch rate. This will in turn release more resources for downstream operators to run, consuming the buffered intermediate data, and bringing the pipeline throughput back to equilibrium.
Ray Data also provides a conservative scheduling policy in which a task is launched only when its output space in the memory can be guaranteed, similar to Figure 4(a). This policy enforces a hard memory limit and never spills, albeit at the risk of under-utilizing of executor slots when memory is limited.
@@ 4.4 Implementation
Current batch and stream processing systems could in principle be modified to use the streaming batch model, but would require fundamental changes to their execution models. For example, Spark would need to support pipelined stage execution, while Flink would need to support zero-downtime reconfiguration. We choose to implement Ray Data on top of Ray because Ray exposes a lower-level execution model based on dynamic task execution. Ray provides powerful features such as automatic data movement, lineage-based recovery [44], and automatic disk spilling [25]. This makes it convenient to build centralized schedulers like Ray Data while leveraging Ray as a decentralized dataplane. However, Ray also treats the task logic, inputs, and outputs as black boxes. Thus, it is difficult to directly extend Ray with data processing-specific features such as dataset partitioning, streaming repartition, and pipeline-aware task scheduling. Instead, we implement Ray Data as a Ray library, which allows us to build such features with minimal changes to the Ray core. Ray Data is written in ~90k Python LoC, including ~45k LoC for the query planner, scheduler and executor logic.
@@ 5 Evaluation
We evaluate a range of heterogeneous workloads with a focus on multimodal batch inference and training. We use benchmarks taken from the MLPerf suite [26, 36], and supplement with additional image-to-image, video-to-video, and video classification benchmarks. We aim to answer:
  • § 5.1: How does the streaming batch model compare to traditional batch or stream processing models when running heterogeneous ML workloads, in terms of throughput and adaptivity vs. fault tolerance?
  • § 5.2: How does distributed and heterogeneous execution improve on throughput per dollar compared to ML-specific single-node data loaders?
  • § 5.3: How well do all systems adapt to memory pressure in heterogeneous settings, and what are the system overheads?
@@
  • Batch processing (Figure 2(a)): Apache Spark 3.5.1.
  • Stream processing (Figure 2(b)): Apache Flink 1.19.0.
  • Single-node stream processing: tf.data [30]. A data loader for ML training. tf.data uses multithreading and can reconfigure the number of threads per operator.
  • Single-node stream processing: PyTorch DataLoader (PyTorch DL) [35]. Similar to above, but uses multiprocessing and requires manual fusing of all operators.
  • Streaming batch: Ray Data (implemented over Ray 2.40.0). We also modify Ray Data to emulate batch processing (Ray Data-staged) and stream processing (Ray Data-static), for apples-to-apples comparison of the execution models in § 2. Ray Data-staged materializes each stage before starting the next, while Ray Data-static sets a static parallelism per operator and disables the adaptive scheduler described in § 4.3.
For training workloads, we compare only against tf.data and PyTorch DataLoader because these systems were custom-built for data preprocessing for training. They do not support distributed execution and target CPU-only preprocessing for a co-located GPU trainer, making it difficult to run batch inference, which alternates between CPU and GPU stages. Thus, we use Spark and Flink as comparisons for batch inference.
(a) Image generation.
(b) Video generation.
@@
(d) Scalability.
Figure 6: (a) Image generation throughput comparison. Ray Data-dynamic (blue) can match the performance of hand-tuned Ray Data-static (red) with static parallelism, and outperform other systems. Ray Data-staged execute synchronously; no results are available until the last stage begins at ~330s. (b) Video generation throughput comparison. Ray Data-dynamic (blue) achieves 28% better throughput compared to Ray Data-static (red) with static parallelism. (c) Fault tolerance comparison. Ray Data can handle isolated executor failures (blue) with negligible impact on throughput, and node failure (red) without restarting the job. The emulated checkpoint-and-restore (pink) leads to job downtime and longer completion time due to recomputation. (d) Scalability analysis. Ray Data is able to achieve strong scaling with respect to the number of nodes.
5.1 Inference: Comparison to distributed batch and stream processing
5.1.1 Image-to-Image Generation
Image-to-image generation uses a generative model to produce a new image from a source image and prompt. Section 5.1.1 consists of the steps: (1) read_images: download the InstructPix2Pix [8] image dataset from S3, (2) decode: decode JPEG images to tensors, (3) preprocess: normalize tensors, (4) Img2ImgModel: use Stable Diffusion [38] to generate a new image based on the input image and prompt, and (5) encode_and_upload: Encode to JPEG and upload to S3.
{listing}
[ht]
ray.data.read_images(INPUT_PATH).map(decode).map(preprocess)
.map_batches(Img2ImgModel, batch_size=B, num_gpus=1)
.map_batches(encode_and_upload, batch_size=B)
Image-to-image generation expressed in Ray Data API.
Figure 6(a) shows throughput over time of the different execution models on 1 g5.2xlarge VM (8 vCPU, 1 A10G GPU). The *-fused baselines fuse all operators. This avoids materialization but limits overall parallelism to the scarcest resource, in this case 1 GPU. The resulting throughput is thus the lowest, confirming that operator fusion is undesirable on heterogeneous resources. Spark-staged and Ray Data-staged disable fusion of CPU and GPU operators to decouple their parallelisms. Stages execute synchronously, so no results are available until the last stage begins at ~330s. The overall throughput is also significantly lower because all intermediate stage results need to be materialized and in this case spilled to disk.
For Flink, we also disable fusion, by setting parallelism to 8 and 1 for CPU and GPU operators, respectively. Flink uses multithreading to share each CPU among its multiple operators (read+decode+preprocess vs. encode_and_upload). Execution is pipelined across physical operators, and only a fraction of intermediate records are materialized at once, so Flink can produce results almost immediately. However, there is significant overhead due to serialization of image data between the Python UDF and the Java-based Flink. Thus, Flink’s throughput is 70% lower than Ray Data’s.
Because of Flink’s serialization overheads, we use Ray Data-static to emulate a stream processing baseline. We manually determine the best static parallelism per operator before execution. This achieves the best throughput, at 4 images/s. Ray Data-dynamic is the default system, with the adaptive scheduler (§ 4.3). This automatically converges to the same throughput as Ray Data-static.
@@ 5.1.1 Image-to-Image Generation
(1) Stream and streaming batch models outperform batch models for heterogeneous pipelines, due to asynchronous stages and lower peak memory footprint, (2) Ray Data’s adaptive scheduler matches the best stream processing baseline (Ray Data-static), with no configuration needed.
Figure 7: Training ResNet-50.
@@
PyTorch DL (stream) 4×4$\times$4 × p4de.24xlarge 2,811 111.3 \$18,192
Ray Data-staged (batch) 4×4$\times$4 × p4de.24xlarge 0, then 4,068 90.3 (-19%) \$14,753 (-19%)
Ray Data (streaming batch) 4×4$\times$4 × p4de.24xlarge 4,075 76.8 (-31%) \$16,275 (-11%)
40×40$\times$40 × g5.2xlarge
Figure 8: Run time and cost for one epoch of Stable Diffusion pre-training.
5.1.2 Video-to-Video Generation
Video-to-video generation uses a generative model for use cases including super-resolution, interpolation, or de-blurring [10, 20, 45]. Compared to § 5.1.1, video generation adds significant memory pressure: each video decodes to many frames. Also, high workload variability is common due to varying video lengths, resolutions, and encodings. Thus, we use this workload to evaluate the system’s ability to adapt compute and memory allocations based on real-time usage.
We run a 3-stage pipeline on a g5.2xlarge VM: (1) download+decode (CPU): Download the YouTube-8M dataset [1] from S3. Video resolution ranges from 320p to 720p, with lengths 2–5 minutes, encoded using H.264. Each video is decoded into multiple 128-frame sequences, the maximum batch size allowed by GPU memory. Earlier videos are low-resolution. (2) generate (GPU): Run the RealBasicVSR [10] super-resolution model to produce HD frames. (3) encode+upload (CPU): Encode the HD frames using H.264, then upload the videos to S3.
Figure 6(b) compares the throughput over time of Ray Data-static and Ray Data-dynamic, the two best systems in § 5.1.1. Ray Data-static emulates stream processing and allocates the pre- and post-processing stages of 4 CPUs each. Throughput is initially high but later drops because this initial configuration is not optimal for the later high-resolution videos, which are more memory-intensive and take longer to download and decode. Meanwhile, with Ray Data-dynamic, the Ray Data scheduler dynamically re-balances tasks according to task duration and memory usage, achieving 28% better throughput.
Takeaway:
Ray Data adapts quickly to changing workloads because tasks can run on any executor and the adaptive scheduler can dynamically reallocate resources.
5.1.3 Fault tolerance in heterogeneous clusters
We first demonstrate Ray Data’s ability to scale with heterogeneous clusters. We run the VideoMAE [41] for video classification, on the Kinetics-700-2020 dataset with 635,000 videos and 871 GB in size. This workload is similar to that in § 5.1.2 except with less postprocessing. Data is stored on Amazon S3. Figure 6(c) shows the throughput over time when processing 10% of the dataset on 1 g5.xlarge node (4 vCPU, 1 GPU). The single-node configuration produces 6.2 videos/s, and is bottlenecked by CPU preprocessing. Ray Data can scale data preprocessing independent of GPU inference. We launch a m7i.2xlarge node (8 vCPU) to create a heterogeneous Ray Data cluster with 12 vCPUs and 1 GPU. Overall throughput then reaches 21.7 videos/s, or 94% of the maximum GPU throughput. The overall job cost is also 60% less, as CPU-only nodes are more cost-efficient. We further test scaling the cluster up to 16 nodes, or 8 GPUs and 96 vCPUs in total. Figure 6(d) shows that Ray Data is able to achieve near-perfect strong scaling.
Fault tolerance.
CPU executor failures are common in heterogeneous clusters mostly due to preprocessing tasks running out of memory. Ray Data can transparently recover from isolated executor failures without interrupting the job. To further evaluate the cost of fault recovery, we intentionally fail the CPU worker node after 10 min, and reconnecting it after another 1 min. We compare Ray Data’s native lineage reconstruction failure recovery against checkpointing, which is commonly used in stream processing systems. Figure 6(c) shows that Ray Data’s throughput drops to that of the remaining node during worker node failure, and restores when the node rejoins without interrupting the job. In comparison, we implement a checkpointing-based recovery, by saving progress every 5 min, then restarting the job to load from the last checkpoint when failure occurs. As expected, the system makes no progress until $t=18\mathrm{min}$, due to having to restart the job and perform redundant work.
@@ 5.1.3 Fault tolerance in heterogeneous clusters
Ray Data enables scaling with heterogeneous clusters to reduce overall cost by 60%, despite using more nodes. Compared to stream processing systems that use global checkpointing, failures at the additional CPU-only nodes have little impact on end-to-end throughput, allowing further cost reduction via spot instances.
5.2 Training: Comparison to ML data loaders
5.2.1 ResNet Training
We run the ResNet-50 ImageNet training benchmark from MLPerf [26]. The data preprocessing pipeline loads images from local disk (local) or cloud storage (S3), decodes, and randomly crops and flips the images. We compare training throughput of tf.data vs. Ray Data on a g5.2xlarge VM. We do not measure PyTorch DataLoader, as tf.data showed comparable or better results for the same benchmark in [30].
Figure 8 shows training throughput over time. tf.data executes data preprocessing using a pool of worker threads running in each GPU trainer process. Thus, the job fate-shares with any preprocessing task that fails due to out-of-memory (OOM). When reading data from local disk, tf.data’s throughput is 19% lower than Ray Data’s because a lower batch size was required to prevent OOM failures. Meanwhile, Ray Data is able to complete because GPU trainer failures are isolated from CPU worker failures, and CPU workers can be respawned in seconds, without impacting pipeline throughput (§ 5.1.3).
When reading data from S3, tf.data is 88% slower than the max GPU throughput because S3 loading is the bottleneck. Meanwhile, Ray Data can use heterogeneous clusters to scale out S3 loading independent of the GPU trainers. By adding a m7i.2xlarge node for data preprocessing, the overall training throughput reaches 93% of the max GPU throughput.

### v5

@@ 4.3.1 Optimistic policy
[…]
3:for $i\leftarrow 1$ to numOpsnumOps do
@@
[…]
The input rate, i.e. the rate at which source tasks are scheduled, must approximate the pipeline’s overall throughput. We use a dynamic memory budget algorithm to regulate the rate at which source tasks are launched. Intuitively, the budget is an optimistic estimate of the memory available for new data partitions to enter the system. When a source task is launched, we deduct its estimated output size from the budget. At every second, the budget is replenished using Algorithm˜2, which estimates the rate at which data leaves the pipeline. For DAGs with multiple sources, we launch each source operator at a rate proportionate to its output size.
@@
  • Consider the first non-source operator: transform. Suppose the current number of available execution slots to run the task is $E_{1}=6$ (out of 8 CPU slots). Assume the average task duration is $T_{1}=12$ seconds. Then the processing time of this stage is $P_{1}=T_{1}/E_{1}\cdot\alpha_{0}=12/6\times 1=2$, where $\alpha_{0}$, the output multiplier, is initialized to 1. In other words, transform takes 2s to process a source partition on average.
@@
If the run-time estimates of each operator’s processing rates are perfectly accurate, i.e. if there is no variance in the processing rates, it can be shown that the schedule produced is optimal. However, when there is variance in processing times or output sizes, the budget algorithm could overestimate the overall processing rate. Nevertheless, the algorithm is stable because it creates a negative feedback loop. If it overestimates the pipeline processing rate, more source tasks might launch and temporarily cause backpressure, or objects in the buffer to spill to disk. However, because these tasks occupy execution slots, they will reduce the parallelism of downstream operators and lower the replenishment rate of the budget, which in turn limits the source task launch rate.
@@ 4.4 Implementation
We implement Ray Data on top of Ray because Ray exposes a low-level execution model based on dynamic task execution, with powerful features such as automatic data movement, lineage-based recovery [41], and disk spilling [24]. This makes it convenient to build centralized schedulers like Ray Data while leveraging Ray as a decentralized dataplane. However, Ray also treats the task logic, inputs, and outputs as black boxes. Thus, it is difficult to directly extend Ray with data processing-specific features such as dataset partitioning and pipeline-aware task scheduling. Instead, we implement Ray Data as a Ray library, which allows us to build such features with minimal changes to the Ray core. Ray Data is an industrial system and is written in ~51k Python LoC, with about 1k, 1k, and 2k LoC for the query planner, scheduler, and map executor logic, respectively.
@@ 5 Evaluation
We evaluate heterogeneous workloads in batch (offline) inference and training, including retrieval-augmented LLM generation (RAG) and multimodal models. We use benchmarks taken from the MLPerf suite [25, 34], and supplement with additional LLM, image, and video benchmarks. We study:
  • § 5.1: How does Ray Data compare to batch or stream processing systems when running heterogeneous ML workloads, in terms of throughput and elasticity?
  • § 5.2: How can distributed and heterogeneous execution reduce cost compared to single-node training data loaders?
  • § 5.3: How well can all systems adapt to memory pressure from intermediate data in heterogeneous settings?
@@
  • Batch processing (Section˜2.2): Apache Spark 3.5.1.
  • Stream processing (Section˜2.3): Apache Flink 1.19.0.
  • Single-node ML data loaders (Section˜2.3): tf.data [28] and PyTorch DataLoader (PyTorch DL) [33]. tf.data can reconfigure the number of threads per operator.
  • Streaming batch: Ray Data (implemented over Ray 2.40.0). We also modify Ray Data to emulate batch processing (Ray Data-staged) and stream processing (Ray Data-static), for apples-to-apples comparison of the execution models. Ray Data-staged materializes each stage before starting the next, while Ray Data-static sets a static parallelism per operator and replaces the dynamic scheduler (§ 4.3) with round-robin partition assignment.
(a) Retrieval-augmented generation.
(b) Video classification.
@@
Figure 7: (a) RAG job completion time. “Staged Batch” runs stages sequentially in one process; its Retrieval run time is the optimal JCT if CPU is the bottleneck. (b) Video classification throughput comparison of batch (Spark and Ray Data-staged), streaming (Flink and Ray Data-static), and streaming batch (Ray Data-dynamic) systems. (c) Fault tolerance comparison for video classification on a heterogeneous cluster with 1 GPU node and 1 CPU-only node. Executor (blue) or node (red) failure is injected at tt=15min. For node failure curves, the node is added back to the cluster at tt=30min. The dashed line is a variant of Ray Data that emulates checkpoint-and-restore. This strategy leads to job downtime after every cluster change and longer completion time. Unmodified Ray Data (solid) smoothly handles executor failure and node failure and addition.
For training workloads, we compare only against tf.data and PyTorch DataLoader because these systems were custom-built for data preprocessing for training. We use Spark, Flink, and Ray Data variants as comparisons for batch inference.
5.1 Inference: Ray Data vs. batch and stream processing
5.1.1 Retrieval-Augmented Generation (RAG)
RAG [23] improves LLM inference accuracy by retrieving relevant documents from a knowledge base to augment the original prompt. Our RAG pipeline consists of three stages: (1) encode (CPU): Encode the input prompt using a pre-trained encoder (Contriever [18]) to produce a dense embedding. (2) retrieve (CPU): Use a FAISS [10] vector index to retrieve the top-$k$ most similar documents from the knowledge base (TriviaQA [19]). (3) generate (GPU): Serve the Llama-3-8B model [14] with vLLM [22] to generate the final output based on the input prompt and retrieved documents. We run the RAG pipeline on 1 node with 8 H200 GPUs and 256 CPUs.
Figure˜7(a) shows the job completion time (JCT) for 100K prompts. We compare Ray Data to a single-process baseline “Staged Batch” that sequentially executes each stage. On 1 GPU, Ray Data achieves 1.32x speedup compared to the baseline. This is because Ray Data runs all stages concurrently and overlaps the CPU- and GPU-based stages, so its overall JCT is similar to that of the generate stage in “Staged Batch”.
We also scale Ray Data to multiple GPUs. On N GPUs, we use Ray Data to set the parallelism of generate to N, creating N vLLM replicas. As shown in Figure˜7(a), compared to 1 GPU, Ray Data is able to achieve 1.88x speedup for 2 GPUs, 3.58x speedup for 4 GPUs, and 6.44x speedup for 8 GPUs. Larger overhead is observed when scaling from 4 to 8 GPUs because the CPU-based stages become the bottleneck.
@@ 5.1.1 Retrieval-Augmented Generation (RAG)
(1) Ray Data outperforms batch processing via asynchronous stage execution, and (2) Ray Data scales LLM inference via data parallelism while saturating the resource bottleneck, either GPU or CPU.
5.1.2 Video Classification
We run the VideoMAE video classification model [38] on the test split of the Kinetics-700-2020 dataset (64,535 videos, 137.3 GB, Amazon S3), on 4 g5.2xlarge nodes (8 vCPUs, 1 NVIDIA A10G GPU each). The operators are: (1) read: download and read binary files from S3, (2) preprocess: decode a video into a series of frames, (3) map_batches(VideoMAE): classify video frames using a GPU-hosted model. Figure˜7(b) shows throughput over time. We include Ray Data-staged and Ray Data-static to emulate batch and streaming systems, respectively.
Batch systems (Spark and Ray Data-staged) execute stages synchronously, so they do not produce results until tt=53min and tt=31min for Spark and Ray Data-staged, respectively. Also, the JCTs are tt=116min and tt=61min respectively because all intermediate stage results are materialized and must be spilled to disk to avoid OOM.
Streaming systems (Flink and Ray Data-static) pipeline across heterogeneous operators so they are able to produce results almost immediately. However, they are also sensitive to the static assignment of operator ranges to executors. Both use a round-robin approach to assign data to executors. Flink’s throughput is more stable than Ray Data-static’s until tt=20min, likely due to finer-grained batching. However, Flink also has significant overheads from serializing and copying data between Java and Python. Thus, we also include Ray Data-static for fair comparison. Ray Data-static’s throughput is unstable due to lack of dynamic load balancing.
[…]
Takeaways:
(1) Streaming and streaming batch systems outperform batch systems through memory-efficient pipelining, (2) Ray Data’s dynamic resource assignment further improves throughput compared to the static assignment often used in streaming systems.
5.1.3 Fault tolerance in heterogeneous clusters
We use the same workload as in Section˜5.1.2 to demonstrate Ray Data’s ability to scale using clusters with heterogeneous node types and to smoothly scale clusters up and down, including during failures. Figure˜7(c) shows the throughput during failure recovery when processing 10% of the dataset using a cluster with 1 g5.xlarge node (4 vCPU, 1 GPU) and 1 m7i.2xlarge node (8 vCPU). For all systems, we inject a node or executor failure at tt=15min. Executor failure kills one worker process. Node failure disconnects the CPU-only node and reconnects it at tt=30min. We compare Ray Data (solid curves) against Ray Data modified to emulate the global checkpoint strategy used in many stream processing systems. The modified Ray Data (dashed curves) takes an empty checkpoint every 6min.
With checkpointing, both executor and node failures force a global rollback, causing downtime until the system recovers the lost work. Downtime under node failure is longer because the system has fewer resources while the CPU-only node is disconnected. Adding the CPU-only node back at tt=30min forces another global rollback. Meanwhile, Ray Data has no noticeable throughput drop under executor failure, and scales smoothly with cluster size during node removal and addition.
Takeaways:
Compared to streaming systems that use global checkpointing, Ray Data achieves similar run-time overhead. Meanwhile, Ray Data also smoothly handles CPU failures and cluster reconfiguration events.
5.2 Training: Ray Data vs. ML data loaders
5.2.1 ResNet Training
We run the ResNet-50 ImageNet training benchmark from MLPerf [25]. The data preprocessing pipeline loads images from local disk (local) or cloud storage (S3), decodes, and randomly crops and flips the images. We compare training throughput vs. tf.data on a g5.2xlarge VM. We do not measure PyTorch DL, as tf.data showed comparable or better results for the same benchmark [28].
Figure˜8(a) shows training throughput over time. tf.data executes data preprocessing using a pool of worker threads running in each GPU trainer process. Thus, the job fate-shares with any preprocessing task that fails due to out-of-memory (OOM). When reading data from local disk, tf.data’s throughput is 19% lower than Ray Data’s because a lower batch size was required to prevent OOM failures. Meanwhile, Ray Data is able to complete because GPU trainer failures are isolated from CPU worker failures, and CPU workers can be respawned in seconds, without impacting pipeline throughput (§ 5.1.3).
When reading data from S3, tf.data is 88% slower than the max GPU throughput because S3 loading is the bottleneck and the GPU’s local CPUs are insufficient. Meanwhile, Ray Data (S3) adds a CPU-only m7i.2xlarge node to scale out S3 loading independent of the GPU trainers. This achieves an overall training throughput of 93% of the maximum GPU throughput.
Takeaways:
Compared to single-node ML data loaders, Ray Data offers: (1) failure isolation between heterogeneous resources, and (2) ability to leverage heterogeneous node types.
5.2.2 Pre-Training Stable Diffusion
Pre-training of large multimodal models is one of the most demanding heterogeneous workloads. We run the Stable Diffusion (SD) pre-training pipeline shown in Figure˜1(b) and compare different execution modes. We execute 1 training epoch over a dataset of 2B images, on a cluster of $4\times$ p4de.4xlarge nodes, with 8 A100 GPUs each. This pipeline requires both CPUs and GPUs for data preprocessing: (1) loadText/loadImage+clip (CPU): Load and preprocess pairs of image and text, (2) Encoder (GPU): Use pre-trained encoder models for image and text to produce embeddings, and (3) UNet.train() (GPU): Train SD, using PyTorch fully-sharded data parallelism (FSDP) [46].
@@ 5.2.2 Pre-Training Stable Diffusion
PyTorch DL (stream) $4\times$ p4de.24xlarge 2,811 111.3 \$18,192
Ray Data-staged (batch) $4\times$ p4de.24xlarge 0, then 4,068 90.3 (-19%) \$14,753 (-19%)
Ray Data (streaming batch) $4\times$ p4de.24xlarge 4,075 76.8 (-31%) \$16,275 (-11%)
$40\times$ g5.2xlarge
Figure 8: (a) Training ResNet-50, loading data from local disk vs. cloud storage (S3). All experiments run on 1 g5.2xlarge node (1 NVIDIA A10G GPU, 8 vCPUs), except for Ray Data (S3), which uses an additional m7i.2xlarge (8 vCPUs) node to scale out S3 loading. (b) Run time and cost for one epoch of Stable Diffusion pre-training.
Figure 9: Run times for systems under different memory limits on a synthetic benchmark. Grey means the system is unable to finish due to OOM. Ray Data(-Part.) means Ray Data without dynamic repartition (Section˜4.2.1). Ray Data(-Adapt.) means Ray Data without adaptive memory-aware scheduling (Section˜4.3.1).
Figure˜8(b) shows training throughput and total cost. PyTorch DL is a data loader custom-built for PyTorch that statically assigns work to a process pool. Its throughput is lowest because Encoder preprocessing competes with trainers for GPU memory. Ray Data-staged emulates batch processing by running data preprocessing as an offline job and storing precomputed embeddings in cloud storage. This is preferable if the same embeddings are used multiple times. Ray Data-staged achieves 19% higher throughput because UNet can use more GPU memory and therefore a larger batch size.
Ray Data runs all data preprocessing concurrently with training. This is preferable if using random transforms or for iterative development. Ray Data also disaggregates Encoder and UNet, placing Encoder onto cheaper A10G GPUs (g5.2xlarge). This results in 31% better throughput than PyTorch DL, because UNet has full GPU resources, and 15% better throughput than Ray Data-staged, because embeddings are kept in memory.
@@
Compared to existing ML data loaders, Ray Data can: (1) be used for both batch and online data preprocessing, and (2) leverage clusters with heterogeneous GPUs.
5.3 Microbenchmarks
5.3.1 Memory-aware pipelining
We evaluate how batch, streaming, and streaming batch systems schedule heterogeneous pipelines with and without memory pressure. The 3-stage pipeline is: (1) Load (CPU): 160 tasks, each producing 500 1 MB rows after 5s, (2) Transform (CPU): sleep for 0.5s per row, then return a new 1 MB row, (3) Inference (GPU): 0.5s per batch of 100 rows. We use 1 m6i.2xlarge node with 8 vCPUs, 4 simulated GPU slots, and 32 GB RAM. The theoretical best job completion time with unlimited memory is $(160\times 5s+800\times 0.5s)/8=150s$.
Figure˜9 shows job completion time vs. total memory limit. We limit memory through system-specific configurations, e.g., executor memory for Spark. We use POSIX rlimit to verify that each system respects its memory limit, and tune each system’s parallelism (e.g., executor count) if not.
Spark materializes all data between stages, achieving at best $2.35\times$ optimal run time. At 12–14GB memory, Spark must use fewer executors resulting in $4.34\times$ the optimal run time, and at lower memory limits, Spark OOMs. This is because Spark uses static partitioning, which is sensitive to memory pressure (§ 2.2).
Flink is less sensitive to the memory limit than Spark and achieves up to $1.68\times$ optimal. This is because executors dynamically batch outputs and apply backpressure to avoid OOM (§ 2.3). Under lower memory limits, Flink must run fewer executors because each CPU slot is multiplexed among multiple physical operator threads. This results in up to $2\times$ worse throughput.
tf.data offers an adaptive scheduler and memory budget, similar to Ray Data, albeit non-distributed. However, we found that the memory budget was not always enforced, requiring manual tuning of the thread count. tf.data achieves the same throughput as Ray Data at 16 GB memory limit, but OOMs at lower memory limits.
Ray Data achieves $1.3\times$ the optimal run time at all memory limits except the lowest. We also conduct ablation studies. Ray Data(-Part.) disables Ray Data’s dynamic repartition (Section˜4.2.1), resulting in too-large partitions similar to Spark. Ray Data(-Adapt.) uses Ray Data’s pessimistic policy instead of its optimistic policy (§ 4.3.1), resulting in 10–88% worse performance than Ray Data. Ray Data is also less sensitive than Flink to memory pressure because the system explicitly time-slices executors, instead of using multithreading.